OpenStudy (anonymous):

Max is in a pie eating contest that lasts 1 hour. Each torte that he eats takes 2 minutes. Each apple pie that he eats takes 3 minutes. He receives 4 points for each torte and 5 points for each pie. What should Max eat so as to get the most points. Could anyone formulate the constraints? Thanks in advance.

OpenStudy (anonymous):

@hba : Could you formulate this one?

OpenStudy (anonymous):

@Murry : Give a try..:)

terenzreignz (terenzreignz):

Think of it this way... How many points would he get in 6 minutes of eating only torte? What about 6 minutes of eating only apple pie?

OpenStudy (anonymous):

Can you write down the equation? I have the equation with me but I'm bit confused

terenzreignz (terenzreignz):

I'm afraid that was all intuitive; I didn't come up with any equation :(

OpenStudy (anonymous):

Oh okaay..Thanks. I know the answer but I'm confused with one equation...:P

OpenStudy (anonymous):

I don't understand the question.. Wish I could help, told you i wasn't smart):

OpenStudy (anonymous):

It's okay JJ. THanks..:)

OpenStudy (anonymous):

Okay(:

OpenStudy (anonymous):

T= number of torte A = number of apple pies N = total numbers of points so equation 1 : 2T + 3A = 60 ( this is the time equation in minutes also the constrain) second equation is : 4T + 5A = N

OpenStudy (anonymous):

@rvelez33 : Thanks. I wrote the something similar Actualy it should be 2T+3A<=60 But in the text it's given as 2T+3A <=24..:(